# Scalable and consistent embedding of probability measures into Hilbert spaces via measure quantization

### v1

@@ 3 Statistical properties of probability measures embeddings via measure quantization
$\min\limits_{a\in(\Sigma_{K})^{N},\ X\in F_{K}}\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}\Bigl{(}\sum\limits_{k=1}^{K}a^{(i)}_{k}\delta_{x_{k}},\mu^{(i)}\Bigr{)},$ (12)
as introduced in [16] for a.c. probability distributions. The following result shows that optimizing (12) is equivalent to $K$-points quantization of the mean measure.
@@
Let $(\mu^{(i)})_{1\leq i\leq N}$ be arbitrary probability measures and let $\overline{\mu}=\frac{1}{N}\sum_{i=1}^{N}\mu^{(i)}$ be the mean measure. Suppose that the cardinality of the support of $\overline{\mu}$ is larger than $K$. Then,
$\min\limits_{a\in(\Sigma_{K})^{N},\ X\in F_{K}}\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}\Bigl{(}\sum\limits_{k=1}^{K}a^{(i)}_{k}\delta_{x_{k}},\mu^{(i)}\Bigr{)}$ $=\min\limits_{X\in{(\mathbb{R}^{d})^{N}}}$ $\frac{1}{N}\sum_{i=1}^{N}W_{2}^{2}\Bigl{(}\sum\limits_{k=1}^{K}\mu^{(i)}(\widetilde{V}_{x_{k}})\delta_{x_{k}},\mu^{(i)}\Bigr{)}$ (13)
$=\min\limits_{X\in(\mathbb{R}^{d})^{N}}$ $W_{2}^{2}\Bigl{(}\sum\limits_{k=1}^{K}\overline{\mu}(\widetilde{V}_{x_{k}})\delta_{x_{k}},\overline{\mu}\Bigr{)}$
For a minimizer $\bar{X}=(\bar{x}_{1},\cdots,\bar{x}_{K})$ of (13), that is a $K$-point quantization of $\overline{\mu}$, we then define the quantized measures for $1\leq i\leq N$ by
$\bar{\nu}^{(i)}_{K}=\sum\limits_{k=1}^{K}\bar{a}^{(i)}_{k}\delta_{\bar{x}_{k}},\mbox{ with }\bar{a}^{(i)}_{k}=\mu^{(i)}(\widetilde{V}_{\bar{x}_{k}}),$ (14)
where $(\widetilde{V}_{\bar{x}_{k}})_{1\leq k\leq K}$ is the Voronoï partition (9) associated to $\bar{X}$. The measure $\bar{\nu}^{(i)}_{K}$ is therefore a discrete probability measure supported on $K$ points that is an approximation of $\mu^{(i)}$ in the sense of the minimization problem (12). The measures $(\bar{\nu}^{(i)}_{K})_{1\leq i\leq N}$ differ in their weights but share the same support $\bar{X}$. In a slight abuse of language, we will refer to $\bar{\nu}^{(i)}_{K}$ as a quantized version of $\mu^{(i)}$, even though it is not the optimal quantization $\tilde{\nu}^{(i)}_{K}$ of $\mu^{(i)}$ given in (11).
@@
Proposition 3.1 remains true without the compactness of $\mathcal{X}$, under the assumption of $2$-order moments of the measures.
@@
$(ii)$ If all the measures $(\mu^{(i)})_{1\leq i\leq N}$ are discrete, then the definition (14) of the probability measure $(\bar{\nu}^{(i)}_{K})_{1\leq i\leq N}$ is specific to the chosen Voronoï partition (9) associated to an enumeration of $\bar{X}$. In this setting, a minimizer $\bar{a}$ of (12) is not necessarily unique, and another enumeration order of $\bar{X}$ may lead to a slightly different set of quantized measures, depending on the intersection between the points clouds and the boundaries of the Voronoï cells.
@@
$V_{k}:=\widetilde{V}_{\bar{x}_{k}},\quad\mbox{for all}\ 1\leq k\leq K.$ (15)
3.1 Consistency of measures quantization
The following result shows the consistency of both quantization methods by leveraging the quantization error function $\varepsilon_{K}(\cdot)$ defined in (10).
Proposition 3.4.
Let $\mathbb{P}^{N}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\mu^{(i)}}$, ¯$\overline{\mathbb{P}}^{N}_{K}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\overline{\nu}^{(i)}_{K}}$ and ~$\widetilde{\mathbb{P}}^{N}_{K}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\tilde{\nu}^{(i)}_{K}}$. Then,
$\mathcal{W}_{2}^{2}(\overline{\mathbb{P}}^{N}_{K},\mathbb{P}^{N})$ $=\frac{1}{N}\sum_{i=1}^{N}W_{2}^{2}(\mu^{(i)},\overline{\nu}^{(i)}_{K})=\varepsilon_{K}(\overline{\mu})$ (16)
@@ 3.1 Consistency of measures quantization
$\mathcal{W}_{2}^{2}(\widetilde{\mathbb{P}}^{N}_{K},\mathbb{P}^{N})$ $=\frac{1}{N}\sum_{i=1}^{N}W_{2}^{2}(\mu^{(i)},\tilde{\nu}^{(i)}_{K})=\frac{1}{N}\sum_{i=1}^{N}\varepsilon_{K}(\mu^{(i)}).$ (17)
@@
Proposition 3.4 and the consistency results (16) and (17) could be extended to more general cost functions $c$ satisfying smoothness conditions such as the so-called $x$-regularity from [23][Definition 1].
In other words, Proposition 3.4 shows the convergence of the empirical measures ¯$\overline{\mathbb{P}}^{N}_{K}$ and ~$\widetilde{\mathbb{P}}^{N}_{K}$ towards $\mathbb{P}^{N}$ at the rate $\mathop{}\mathopen{}O\mathopen{}\left(K^{-2/d}\right)$ as $K$ goes to +$\infty$+ ∞.
@@
By definition of optimal quantization (6), one has the inequality $W_{2}^{2}(\mu^{(i)},\tilde{\nu}^{(i)}_{K})\leq W_{2}^{2}(\mu^{(i)},\overline{\nu}^{(i)}_{K})$ for all $1\leq i\leq N$. Therefore, we deduce from Proposition 3.4 that $\mathcal{W}_{2}^{2}(\widetilde{\mathbb{P}}^{N}_{K},\mathbb{P}^{N})\leq\mathcal{W}_{2}^{2}(\overline{\mathbb{P}}^{N}_{K},\mathbb{P}^{N})$. Hence, ~$\widetilde{\mathbb{P}}^{N}_{K}$ is a better approximation of $\mathbb{P}^{N}$ than ¯$\overline{\mathbb{P}}^{N}_{K}$. Still, the rates of convergence of $\mathcal{W}_{2}^{2}(\widetilde{\mathbb{P}}^{N}_{K},\mathbb{P}^{N})$ and $\mathcal{W}_{2}^{2}(\overline{\mathbb{P}}^{N}_{K},\mathbb{P}^{N})$ are both scaling as $O(K^{-2/d})$. Moreover, when the $\mu^{(i)}$’s are discrete, mean-measure quantization has computational advantages over optimal quantization of each input measure for moderate to large values of $N$ (see Section 4.1).
@@
We simplify notation by denoting $\tilde{\nu}^{(i)}_{K}$ (resp. ~$\widetilde{\mathbb{P}}^{N}_{K}$), defined in (11), and $\bar{\nu}^{(i)}_{K}$ (resp. ¯$\overline{\mathbb{P}}^{N}_{K}$) defined in (14), by $\nu^{(i)}_{K}$ (resp. $\mathbb{P}^{N}_{K}$). We also write $\varepsilon_{K}=\mathcal{W}_{2}^{2}(\mathbb{P}_{K}^{N},\mathbb{P}^{N})$, where $\varepsilon_{K}=\frac{1}{N}\sum_{i=1}^{N}\varepsilon_{K}(\mu^{(i)})$ in the case of optimal quantization of each input measure, and $\varepsilon_{K}=\varepsilon_{K}(\overline{\mu})$ in the case of mean-measure quantization.
Since $\mathcal{X}$ is a compact set, so is the metric space $(\mathcal{P}(\mathcal{X}),W_{2})$ [44][Remark 6.17]. Then, by [40][Theorem 5.9], $\mathcal{W}_{2}(\mathbb{P}^{N}_{K},\mathbb{P})\rightarrow 0$ if and only if $\mathbb{P}^{N}_{K}\rightarrow\mathbb{P}^{N}$ in the sense of weak convergence of distributions, or in other words for any bounded continuous function $f:\mathcal{P}(\mathcal{X})\rightarrow\mathbb{R}$, it holds that $\int f(\nu)\mathrm{d}\mathbb{P}^{N}_{K}(\nu)\overset{K\to+\infty}{\longrightarrow}\int f(\mu)\mathrm{d}\mathbb{P}^{N}(\mu)$. Proposition 3.4 implies the consistency of numerous statistics computed from the quantized measures $(\nu^{(i)}_{K})_{1\leq i\leq N}$ and also allows to show convergence in the MMD sense.
@@
One has
$\frac{1}{N}\sum_{i=1}^{N}\mathrm{MMD}^{2}(\mu^{(i)},\nu^{(i)}_{K})\leq\varepsilon_{K}\overset{K\rightarrow\infty}{\longrightarrow}0.$ (18)
3.2 Statistics from the quantized measures
We focus here on how statistics computed from the quantized measures $(\nu^{(i)}_{K})_{i=1}^{N}$, defined either by (11) or (14), relate to statistics from the input measures $(\mu^{(i)})_{i=1}^{N}$.
3.2.1 Wasserstein barycenter
@@
3.2.2 Statistical dispersion
For a set of measures $\mu=(\mu^{(i)})_{i=1}^{N}$, we define its dispersion as the sum of squares $\mathrm{SS}(\mu)=\frac{1}{N^{2}}\sum_{i,j=1}^{N}W_{2}^{2}(\mu^{(i)},\mu^{(j)}).$ The following shows that $\mathrm{SS}(\nu_{K})$, for $\nu_{K}=(\nu^{(i)}_{K})_{i=1}^{N}$, is controlled by $\mathrm{SS}(\mu)$ and the quantization error $\varepsilon_{K}$ defined in Remark 3.7.
@@ 3.2.2 Statistical dispersion
Guaranteeing that the pairwise distance $W_{2}(\nu^{(i)}_{K},\nu^{(j)}_{K})$ is a good approximation of $W_{2}(\mu^{(i)},\mu^{(j)})$ is essential as many machine learning tasks rely on comparing pairs of data. For mean-measure quantization (12), we provide below a result on pairwise distances when the input measures are all a.c., which only depends on the quantization.
@@
Suppose that the probability measures $(\mu^{(i)})_{i=1}^{N}$ are a.c. Then, one has
@@
with $\mathrm{diam}(V_{k})=\max\limits_{x,y\in V_{k}}\|x-y\|^{2}$ and $(V_{k})_{k=1}^{K}$ the Voronoï cells (15) obtained from the $K$-points quantization of $\bar{\mu}$.

### v3

@@ 3.1 Two quantization schemes
$\min\limits_{a\in(\Sigma_{K})^{N},\ X\in F_{K}}\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}\Bigl{(}\sum\limits_{k=1}^{K}a^{(i)}_{k}\delta_{x_{k}},\mu^{(i)}\Bigr{)},$ (3.2)
as introduced in [20] for a.c. probability distributions. The following result shows that optimizing (3.2) is equivalent to $K$-points quantization of the mean measure.
@@
Let $(\mu^{(i)})_{1\leq i\leq N}$ be arbitrary probability measures with support included in a compact set $\mathcal{X}\subset\mathbb{R}^{d}$ and let $\overline{\mu}=\frac{1}{N}\sum_{i=1}^{N}\mu^{(i)}$ be the mean measure. Suppose that the cardinality of the support of $\overline{\mu}$ is larger than $K$. Then,
$\min\limits_{a\in(\Sigma_{K})^{N},\ X\in F_{K}}\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}\Bigl{(}\sum\limits_{k=1}^{K}a^{(i)}_{k}\delta_{x_{k}},\mu^{(i)}\Bigr{)}$ $=\min\limits_{X\in{(\mathbb{R}^{d})^{K}}}$ $\frac{1}{N}\sum_{i=1}^{N}W_{2}^{2}\Bigl{(}\sum\limits_{k=1}^{K}\mu^{(i)}(U_{x_{k}})\delta_{x_{k}},\mu^{(i)}\Bigr{)}$ (3.3)
$=\min\limits_{X\in(\mathbb{R}^{d})^{K}}$ $W_{2}^{2}\Bigl{(}\sum\limits_{k=1}^{K}\overline{\mu}(U_{x_{k}})\delta_{x_{k}},\overline{\mu}\Bigr{)}$
For a minimizer $\bar{X}=(\bar{x}_{1},\cdots,\bar{x}_{K})$ of (3.3), that is a $K$-point quantization of $\overline{\mu}$, we then define the quantized measures for $1\leq i\leq N$ by
$\bar{\nu}^{(i)}_{K}=\sum\limits_{k=1}^{K}\bar{a}^{(i)}_{k}\delta_{\bar{x}_{k}},\mbox{ with }\bar{a}^{(i)}_{k}=\mu^{(i)}(U_{\bar{x}_{k}}),$ (3.4)
where $(U_{\bar{x}_{k}})_{1\leq k\leq K}$ is the Voronoï partition (2.10) associated to $\bar{X}$. The measure $\bar{\nu}^{(i)}_{K}$ is therefore a discrete probability measure supported on $K$ points that is an approximation of $\mu^{(i)}$ in the sense of the minimization problem (3.2). The measures $(\bar{\nu}^{(i)}_{K})_{1\leq i\leq N}$ differ in their weights but share the same support $\bar{X}$. In a slight abuse of language, we will refer to $\bar{\nu}^{(i)}_{K}$ as a quantized version of $\mu^{(i)}$, even though it is not the optimal quantization $\tilde{\nu}^{(i)}_{K}$ of $\mu^{(i)}$ given in (3.1).
@@
Proposition 3.1 remains true without the compactness assumption on $\mathcal{X}$, under finite $2$-order moments of the measures.
@@
$(ii)$ If all the measures $(\mu^{(i)})_{1\leq i\leq N}$ are discrete, then the definition (3.4) of the probability measure $(\bar{\nu}^{(i)}_{K})_{1\leq i\leq N}$ is specific to the chosen Voronoï partition (2.10) associated to an enumeration of $\bar{X}$. In this setting, a minimizer $\bar{a}$ of (3.2) is not necessarily unique, and another enumeration order of $\bar{X}$ may lead to a slightly different set of quantized measures, depending on the intersection between the points clouds and the boundaries of the Voronoï cells.
@@
$V_{k}:=U_{\bar{x}_{k}},\quad\mbox{for all}\ 1\leq k\leq K.$ (3.5)
3.2 Main result
The following result shows the consistency of both quantization methods by leveraging the quantization error function $\varepsilon_{K}(\cdot)$ defined in (2.11).
Theorem 3.4.
[…]
$\mathcal{W}_{2}^{2}(\overline{\mathbb{P}}^{N}_{K},\mathbb{P}^{N})$ $=\frac{1}{N}\sum_{i=1}^{N}W_{2}^{2}(\mu^{(i)},\overline{\nu}^{(i)}_{K})=\varepsilon_{K}(\overline{\mu})$ (3.6)
@@ 3.2 Main result
$\mathcal{W}_{2}^{2}(\widetilde{\mathbb{P}}^{N}_{K},\mathbb{P}^{N})$ $=\frac{1}{N}\sum_{i=1}^{N}W_{2}^{2}(\mu^{(i)},\tilde{\nu}^{(i)}_{K})=\frac{1}{N}\sum_{i=1}^{N}\varepsilon_{K}(\mu^{(i)}).$ (3.7)
@@
Theorem 3.4 and the consistency results (3.6) and (3.7) could be extended for more general cost functions than the squared Euclidean cost in the $2$-Wasserstein distance (2.2). For example, for a smooth cost verifying the so-called $x$-regularity from [26][Definition 1], the results would hold.
In other words, Theorem 3.4 shows the convergence of the empirical measures ¯$\overline{\mathbb{P}}^{N}_{K}$ and ~$\widetilde{\mathbb{P}}^{N}_{K}$ towards $\mathbb{P}^{N}$ at the rate $\mathop{}\mathopen{}O\mathopen{}\left(K^{-2/d}\right)$ as $K$ goes to +$\infty$+ ∞.
@@
By definition of optimal quantization (2.6), one has the inequality $W_{2}^{2}(\mu^{(i)},\tilde{\nu}^{(i)}_{K})\leq W_{2}^{2}(\mu^{(i)},\overline{\nu}^{(i)}_{K})$ for all $1\leq i\leq N$. Therefore, we deduce from Theorem 3.4 that $\mathcal{W}_{2}^{2}(\widetilde{\mathbb{P}}^{N}_{K},\mathbb{P}^{N})\leq\mathcal{W}_{2}^{2}(\overline{\mathbb{P}}^{N}_{K},\mathbb{P}^{N})$. Hence, ~$\widetilde{\mathbb{P}}^{N}_{K}$ is a better approximation of $\mathbb{P}^{N}$ than ¯$\overline{\mathbb{P}}^{N}_{K}$. Still, the rates of convergence of $\mathcal{W}_{2}^{2}(\widetilde{\mathbb{P}}^{N}_{K},\mathbb{P}^{N})$ and $\mathcal{W}_{2}^{2}(\overline{\mathbb{P}}^{N}_{K},\mathbb{P}^{N})$ are both scaling as $O(K^{-2/d})$. Moreover, when the $\mu^{(i)}$’s are discrete, mean-measure quantization has computational advantages over optimal quantization of each input measure for moderate to large values of $N$ (see Section 5.1).
@@
In the following, as both quantization methods exhibit similar behavior, we simplify notation by denoting both $\tilde{\nu}^{(i)}_{K}$ (resp. ~$\widetilde{\mathbb{P}}^{N}_{K}$), defined in (3.1), and $\bar{\nu}^{(i)}_{K}$ (resp. ¯$\overline{\mathbb{P}}^{N}_{K}$) defined in (3.4), with a single notation $\nu^{(i)}_{K}$ (resp. $\mathbb{P}^{N}_{K}$). We also write $\varepsilon_{K}=\mathcal{W}_{2}^{2}(\mathbb{P}_{K}^{N},\mathbb{P}^{N})$, where $\varepsilon_{K}=\frac{1}{N}\sum_{i=1}^{N}\varepsilon_{K}(\mu^{(i)})$ in the case of optimal quantization of each input measure (see (3.7)), and $\varepsilon_{K}=\varepsilon_{K}(\overline{\mu})$ in the case of mean-measure quantization (see (3.6)).
Since $\mathcal{X}$ is a compact set, so is the metric space $(\mathcal{P}(\mathcal{X}),W_{2})$ [51][Remark 6.17]. Then, by [45][Theorem 5.9], $\mathcal{W}_{2}(\mathbb{P}^{N}_{K},\mathbb{P})\rightarrow 0$ if and only if $\mathbb{P}^{N}_{K}\rightarrow\mathbb{P}^{N}$ in the sense of weak convergence of distributions, or in other words for any bounded continuous function $f:\mathcal{P}(\mathcal{X})\rightarrow\mathbb{R}$, it holds that $\int f(\nu)\mathrm{d}\mathbb{P}^{N}_{K}(\nu)\overset{K\to+\infty}{\longrightarrow}\int f(\mu)\mathrm{d}\mathbb{P}^{N}(\mu)$. Therefore, one can deduce from Theorem 3.4 the consistency of numerous statistics computed from the quantized measures $(\nu^{(i)}_{K})_{1\leq i\leq N}$ as well as convergence in the MMD sense.
@@
Let $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ be a positive definite kernel. For any probability measures $\mu,\nu\in\mathcal{P}(\mathcal{X})$, we write $\mathrm{MMD}_{k}(\mu,\nu)=\|\phi(\mu)-\phi(\nu)\|_{\mathcal{H}}$, where $\phi$ and $\mathcal{H}$ are respectively the kernel mean embedding defined in (2.5) and then RKHS associated to $k$. Assume that there exists a constant $C>0$ such that, for all $x,y\in\mathcal{X}$,
$k(x,x)+k(y,y)-2k(x,y)\leq C^{2}\|x-y\|^{2}.$ (3.8)
Then, one has that
$\frac{1}{N}\sum_{i=1}^{N}\mathrm{MMD}_{k}^{2}(\mu^{(i)},\nu^{(i)}_{K})\leq C^{2}\varepsilon_{K}\overset{K\rightarrow\infty}{\longrightarrow}0.$
Corollary 3 in [50] gives mild assumptions for condition (3.8) to hold. It is in particular true for the popular Gaussian kernel.
3.3 Consistent estimation from empirical measures
For an arbitrary measure $\mu$, let us denote $\hat{\mu}=\frac{1}{m}\sum_{j=1}^{m}\delta_{x_{j}}$ the usual empirical measure for $x_{1},\ldots,x_{m}\overset{i.i.d.}{\sim}\mu$. In the above section, we assumed access to the true target measures $(\mu^{(i)})_{i=1}^{N}$. However, in practice when $\mu^{(i)}$ is a.c., we only have access to $\mu^{(i)}$ via random samples, and we can thus only perform quantization from the empirical measure $\hat{\mu}^{(i)}$. Let us denote ^$\hat{\mathbb{P}}_{K}^{N}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\hat{\nu}^{(i)}_{K}}$, where each $\hat{\nu}^{(i)}_{K}$ is obtained by optimal quantization of the input measure $\hat{\mu}^{(i)}$ as described in equation (3.1) in Section 3.1. The following result shows convergence in high probability of ^$\hat{\mathbb{P}}_{K}^{N}$ towards $\mathbb{P}^{N}$.
Theorem 3.9.
Let $(\mu^{(i)})_{1\leq i\leq N}$ be $N$ probability measures supported on a compact set $\mathcal{X}\subset\mathbb{R}^{d}$ with absolutely continuous parts $f^{(i)}\neq 0$. For $\delta>0$, all sufficiently large values of $(m_{i})_{1\leq i\leq N}$ and $K=\max\limits_{1\leq i\leq N}C\cdot m(f^{(i)})\cdot m_{i}^{d/(2d+4)}$, it holds that:
$\mathcal{W}_{2}^{2}(\hat{\mathbb{P}}^{N}_{K},\mathbb{P}^{N})\leq C^{2}\frac{\delta^{2}+\log(N)}{N}\sum\limits_{i=1}^{N}m(f^{(i)})^{2}\cdot m_{i}^{-1/(d+2)}~{}\text{with probability}~{}1-e^{-\delta^{2}},$
where $C$ is a constant that only depends on the dimension $d$ and $m(f)=\int_{\mathcal{X}}f(x)^{d/(d+2)}\mathrm{d}\lambda_{\mathcal{X}}(x)$ with $\lambda_{\mathcal{X}}$ the Lebesgue measure on $\mathcal{X}$.
Note that Theorem 3.9 only holds for quantized measures obtained from the process of optimal quantization of each input measure. Indeed, the arguments used in the proof of Theorem 3.9 cannot be applied when the quantized measures are obtained by mean-measure quantization.
4 Statistics from the quantized measures
We focus here on how statistics computed from the quantized measures $(\nu^{(i)}_{K})_{i=1}^{N}$, defined either by (3.1) or (3.4), relate to statistics computed from the input measures $(\mu^{(i)})_{i=1}^{N}$.
4.1 Wasserstein barycenter
@@
4.2 Statistical dispersion
For a set of measures $\mu=(\mu^{(i)})_{i=1}^{N}$, we define its dispersion as the sum of squares $\mathrm{SS}(\mu)=\frac{1}{N^{2}}\sum_{i,j=1}^{N}W_{2}^{2}(\mu^{(i)},\mu^{(j)}).$ The following shows that $\mathrm{SS}(\nu_{K})$, for $\nu_{K}:=(\nu^{(i)}_{K})_{i=1}^{N}$, is controlled by $\mathrm{SS}(\mu)$ and the quantization error $\varepsilon_{K}$ defined in Remark 3.7.
@@ 4.2 Statistical dispersion
Guaranteeing that the pairwise distance $W_{2}(\nu^{(i)}_{K},\nu^{(j)}_{K})$ is a good approximation of $W_{2}(\mu^{(i)},\mu^{(j)})$ is essential as many machine learning tasks rely on comparing pairs of data. For mean-measure quantization (3.2), we provide below a result on pairwise distances when the input measures are all a.c.
@@
Suppose that the probability measures $(\mu^{(i)})_{i=1}^{N}$ are a.c. Then, one has for $1\leq i,j\leq N$ and the mean-measure quantization approach in equation (3.4)
@@
with $\mathrm{diam}(V_{k})=\max\limits_{x,y\in V_{k}}\|x-y\|^{2}$ and $(V_{k})_{k=1}^{K}$ the Voronoï cells (3.5) obtained from the $K$-points quantization of $\bar{\mu}$.